Solve each problem and provide the complete solution.
note: topic is all about Gravitational force

Two objects, A and B, have weights of 5kg and 6kg respectively. If they are 2m away from each other, what is the force of attraction between them?
1
Expert's answer
2020-11-27T00:40:24-0500

We can find the force of attraction between the object A and B from the Newton’s law of universal gravitation:


"F=G\dfrac{m_Am_B}{r_{AB}^2},""F=6.672\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\dfrac{5\ kg\cdot 6\ kg}{(2\ m)^2}=5\cdot10^{-10}\ N."

Answer:

"F=5\cdot10^{-10}\ N."